OpenStudy (gabylovesyou):

Which polynomial identity will prove that 117 = 125 − 8? Difference of Squares Difference of Cubes Sum of Cubes Square of a Binomial

OpenStudy (gabylovesyou):

@hartnn @ganeshie8

OpenStudy (bahrom7893):

125=5^3

OpenStudy (bahrom7893):

8=2^3

RhondaSommer (rhondasommer):

a³ - b³ = (a - b)(a² + a b + b²) 125 - 8 = 5³ - 2³ = (5 - 2)(5² + 5·2 + 2²) = 3·39 = 117

OpenStudy (bahrom7893):

What do you think?

OpenStudy (gabylovesyou):

difference of cubes ?

RhondaSommer (rhondasommer):

yup :)

OpenStudy (damonte212):

difference of cubes a³ - b³ = (a - b)(a² + a b + b²) 125 - 8 = 5³ - 2³ = (5 - 2)(5² + 5·2 + 2²) = 3·39 = 117

OpenStudy (gabylovesyou):

Thank you!! :)